OpenStudy (anonymous):

Find A1? Given: Sn: -342, An: -512, R: -2

OpenStudy (princeharryyy):

what is Sn and An

OpenStudy (princeharryyy):

@Jillian_Summer

OpenStudy (anonymous):

\[A _{n}=A _{1}R ^{n-1}\]\[S _{n}=A _{1}(R ^{n}-1)/(R-1)\] solve and get the answer

OpenStudy (anonymous):

I think your second equation is off @sumitprataps : \[S _{n}= A _{1}(1-R ^{n})/(1-R)\] I have the formulas and know how to plug in the information given to me; however solving it has proven challenging to both myself and professors. If you had a step by step of how to solve it I would really appreciate it! :)

OpenStudy (princeharryyy):

both are valid depending on the value of R

OpenStudy (anonymous):

That may be true but for this case you would need to stick with the original formula because if we used R-1 it would give you -3 but using 1-R gives you +3 and the signs (+/-) matter for this type of equation. @princeharryyy

OpenStudy (princeharryyy):

use watever u like the answer will not change but just the theory willl.... @Jillian_Summer

OpenStudy (princeharryyy):

u have to use 1-r according to the theory

OpenStudy (princeharryyy):

they say whenever r is -ve u got to use 1-r

OpenStudy (princeharryyy):

see (1-r^n)/(1-r) = (r^n -1)/(r-1) on taking - common

OpenStudy (princeharryyy):

they both cancels out

OpenStudy (princeharryyy):

from numerator n d denominator

OpenStudy (triciaal):